While standing at the edge of the roof of a building, a man throws a stone upward with an initial speed of 6.43 m/s. The stone subsequently falls to the ground, which is 11.7 m below the point where the stone leaves his hand. At what speed does the stone impact the ground? Ignore air resistance and use g = 9.81 m/s2 for the acceleration due to gravity. impact speed: m/s How much time is the stone in the air? elapsed time:More people end up in U.S. emergency rooms because of fall-related injuries than from any other cause. At what speed u would someone hit the ground who accidentally steps off the top rung of a 9.00 ft tall stepladder? (That step is usually embossed with the phrase Warning! Do not stand on this step.) Ignore the effects of air resistance. Let the +y direction point vertically upward and the origin be on the ground. m/s
Answer #1

1) By using,

v^2 = u^2 + 2*a*s

v = sqrt(6.43^2 + 2*(-9.8)*11.7)

= 13.71 m/s

Now,

let the elapsed time be t,

v = u + at

t = (v - u)/a

= (13.71 - 6.43)/9.8

= 0.74 s